### v1

# From XAI to Stories: A Factorial Study of LLM-Generated Explanation Quality

Abstract
[…]
From XAI to Stories: A Factorial Study of LLM-Generated Explanation Quality
Fabian Lukassen1, Jan Herrmann2, Christoph Weisser1, Benjamin Saefken1,3, Thomas Kneib1 1University of Göttingen  2BASF SE  3TU Clausthal fabian.lukassen@stud.uni-goettingen.de, jan.herrmann@basf.com, christoph.weisser@oxon.org benjamin.saefken@tu-clausthal.de, tkneib@uni-goettingen.de
Figure 1: Overview of our Prediction-to-NLE pipeline: Models generate predictions, XAI methods (SHAP/LIME) compute feature attributions, and LLMs transform these numerical outputs into narratives accessible to lay users.
1 Introduction
The deployment of ML models in critical domains such as healthcare, finance, and energy management has intensified the need for model interpretability (guidotti2018survey). XAI methods like SHAP (lundberg2017unified) and LIME (ribeiro2016why) have emerged as essential tools, providing quantitative feature attributions that reveal which inputs drive model predictions. However, these technical outputs – typically numerical importance scores or attribution values – remain inaccessible to non-expert users who lack the statistical background to interpret them (zytek2024llms).
Consider a model predicting next week’s household energy consumption. A typical SHAP output might show: “lag_1: +0.23, weekofyear: -0.15, holiday_count: +0.08.” While technically precise, a homeowner cannot readily understand that last week’s high usage is pushing the forecast up, that the current season typically has lower demand, or that an upcoming holiday slightly increases expected consumption. This interpretability gap limits the practical impact of XAI methods and hinders the democratization of model transparency.
Recent work has demonstrated that LLMs can transform XAI output into NLEs that non-experts can better understand (Figure 1), with user studies validating this approach (zytek2024explingo; cedro2024graphxain).
Figure 2: Factorial experimental design with four factors and their levels (top), and five G-Eval evaluation criteria (bottom).
However, previous work on transforming XAI outputs into NLEs does not systematically study the effects of ML model choice, XAI method, LLM, and prompting strategy on NLE quality. We address this gap with a factorial study grounded in Design of Experiments (DOE) principles (collins2014factorial), focusing on time-series forecasting. A domain where temporal dependencies and seasonal patterns make XAI outputs especially challenging to interpret. The factorial combination of these four factors (see Figure 2 for their levels) yields 660 NLEs, which we evaluate using G-Eval (liu2023geval), an LLM-as-judge framework with dual judges and four evaluation criteria. Our main contributions are:
  1. 1.
    A factorial study design with statistical analysis (ANOVA, effect sizes, robustness checks) to systematically quantify what factors matter for XAI-to-NLE quality, including a crucial no-XAI baseline absent from prior work.
  2. 2.
    The observation of an “interpretability paradox”: in our setting, SARIMAX yielded lower NLE quality than ML models despite higher prediction accuracy and inherent interpretability.
  3. 3.
[…]
  4. 4.
    Evidence that LLM choice dominates: LLM selection explains far more variance in NLE quality than XAI method, prompting strategy, or ML model choice.
2 Related Work
2.1 Explainable AI Methods
Among post-hoc explanation methods, SHAP and LIME have emerged as the two most widely adopted techniques for interpreting black-box models (garreau2020explaining).
SHAP (lundberg2017unified) provides theoretically grounded attributions based on Shapley values from cooperative game theory. For a prediction $f(x)$ with feature set $\mathcal{F}$, SHAP assigns each feature $i$ an importance value $\phi_{i}$ computed as a weighted average of marginal contributions across all possible feature coalitions $S$:
$\phi_{i}=\sum_{S\subseteq\mathcal{F}\setminus\{i\}}w_{|S|}\left[f(x_{S\cup\{i\}})-f(x_{S})\right]$ (1)
where $w_{|S|}=\frac{|S|!(|\mathcal{F}|-|S|-1)!}{|\mathcal{F}|!}$. The explanation satisfies local accuracy: $f(x)=\phi_{0}+\sum_{i=1}^{M}\phi_{i}$, with $\phi_{0}$ being the baseline prediction.
LIME (ribeiro2016why) takes a different approach, fitting interpretable local surrogate models. Given instance $x$ and black-box model $f$, LIME learns a locally faithful linear model $g$ by solving:
$\xi(x)=\arg\min_{g\in\mathcal{G}}\mathcal{L}(f,g,\pi_{x})+\Omega(g)$ (2)
where $\pi_{x}$ is a proximity measure defining the local neighborhood and $\Omega(g)$ penalizes complexity.
Both methods have been widely adopted (guidotti2018survey), applied in complex machine learning pipelines (semnani2024amachine; semnani2025from), and extended to specialized variants including TreeSHAP (lundberg2020local). However, despite their mathematical rigor, XAI outputs require statistical expertise to interpret, particularly for time-series data where temporal dependencies add complexity (theissler2022explainable). zytek2024llms identify this as a fundamental barrier, noting that “feature importance scores are only useful if users can understand what they mean.”
In contrast, classical statistical models like SARIMAX offer inherent interpretability through coefficients and confidence intervals (box2015time).
2.2 Natural Language Explanations from XAI
The transformation of XAI outputs to natural language has progressed from template-based systems to LLM-based approaches. martens2024xaistories introduced XAIstories, generating narratives from SHAP values, with user studies showing participants found them convincing and useful for communicating with laypeople. zytek2024explingo developed Explingo, also using SHAP, combining XAI-to-narrative transformation with an LLM-based grader. cedro2024graphxain extended this to graph neural networks using GNNExplainer. For time-series specifically, shajalal2024forecastexplainer used DeepLIFT for LSTM-based energy forecasting, while aksu2024xforecast proposed XForecast with simulatability metrics, finding that an LLM’s numerical reasoning ability matters more than model size for accurate explanations. Domain applications include chemistry (wellawatte2023extracting).
2.3 LLM Prompting Strategies
The effectiveness of LLMs depends heavily on prompting. Foundational techniques include zero-shot and few-shot in-context learning (brown2020language), while more advanced strategies target reasoning and reliability. Chain-of-thought (CoT) prompting elicits step-by-step reasoning (wei2022chain), which self-consistency extends by sampling multiple reasoning paths and selecting the most common answer (wang2023selfconsistency). Other approaches include reflexion, where models iteratively critique and refine their outputs (shinn2023reflexion); meta-prompting, which uses structured templates for multi-step reasoning (suzgun2024metaprompting); and role-based prompting, which assigns personas to guide generation (kong2024betterzeroshotreasoning).
2.4 LLM-as-Judge Evaluation
Traditional metrics (BLEU, ROUGE) correlate poorly with human judgment for open-ended text generation (liu2023geval). The LLM-as-judge paradigm has emerged as a scalable alternative (zheng2023judging), with three main evaluation paradigms (gu2024llmasjudge): pointwise (scoring individual outputs independently), pairwise (comparing two outputs to select the better one), and listwise (ranking multiple outputs simultaneously).
liu2023geval introduced G-Eval, a pointwise framework combining chain-of-thought prompting with probability-weighted scoring. G-Eval first generates evaluation steps from a criterion (e.g., “coherence”), then uses these steps to assess outputs via a form-filling paradigm. To obtain fine-grained scores, G-Eval computes:
$\text{Score}_{\text{G-Eval}}(y)=\sum_{i=1}^{N}i\cdot P(\text{score}=i|y,c,d)$ (3)
where $P(\text{score}=i)$ is derived from the softmax over token log-probabilities for score tokens “1” through “NN”, $y$ is the generated text, $c$ is the evaluation context, and $d$ is the dimension-specific rubric. This yields continuous scores (e.g., 3.7) rather than discrete integers, achieving 0.514 Spearman correlation with human judgments.
However, LLM judges exhibit systematic biases (gu2024llmasjudge): positional bias (preference for first or last positions in ranking tasks), self-preference bias (favoring outputs from the same LLM family), and verbosity bias (favoring longer responses). Mitigation strategies include randomizing presentation order, using judges from different model families than generators, and supplementing absolute scores with ranking-based evaluation.
2.5 Positioning Our Work
While previous work demonstrates XAI-to-NLE feasibility, none systematically investigates which factors matter. We fill this gap with a factorial design that varies ML models (including classical vs. ML comparison), XAI methods (including no-XAI baseline), LLMs, and prompting strategies. This enables us to address the following research questions (RQs):
  • RQ1 (XAI Value): Does including XAI outputs improve NLE quality compared to no-XAI baselines, and do SHAP and LIME differ in their effectiveness?
  • RQ2 (LLM Characteristics): Does the choice of the LLM (varying in scale and reasoning capabilities) affect explanation quality?
  • RQ3 (Model Effects): Do different ML models and their predictive quality affect NLE quality?
  • RQ4 (Model Type): Do classical time-series models yield different explanation quality than black-box ML models?
  • RQ5 (Prompting): Which prompting strategy is most effective?
3 Methodology
3.1 Problem Formulation
Algorithm 1 outlines our Prediction-to-evaluated-NLE pipeline.
Algorithm 1 Experimental Pipeline
1:Dataset $\mathcal{D}$, Models $\mathcal{M}$, XAI $\mathcal{X}$, LLMs $\mathcal{L}$, Strategies $\mathcal{S}$
2:Evaluated explanation corpus $\mathcal{E}$
3:Preprocess $\mathcal{D}$; split train/test chronologically
4:// Phase 1: Generation
5:for all $m\in\mathcal{M}$ do
6:  Train $m$; compute $\hat{\mathbf{y}}$ and metrics $(R^{2},\text{MAE},\text{RMSE})$
7:  for all $\chi\in\mathcal{X}$ do
8:   Compute attributions $\boldsymbol{\phi}\leftarrow\chi(m,\mathbf{z})$
9:   for all $\ell\in\mathcal{L}$, $s\in\mathcal{S}$ do
10:     $E\leftarrow\text{GenerateNLE}(\ell,s,\hat{y},\boldsymbol{\phi},\mathbf{c})$
11:     Store $(m,\chi,\ell,s,E)$ in corpus $\mathcal{C}$
12:   end for
13:  end for
14:end for
15:// Phase 2: Evaluation
16:for all $(m,\chi,\ell,s,E)\in\mathcal{C}$ do
17:  $\mathbf{q}\leftarrow\text{Evaluate}(E)$ $\triangleright$ G-Eval scoring
18:  Store $(m,\chi,\ell,s,E,\mathbf{q})$ in $\mathcal{E}$
19:end for
20:return $\mathcal{E}$
Let $\mathbf{X}=\{x_{t}\}_{t=1}^{n}$ be a univariate time series. Given a forecasting model $f$ trained on features $\mathbf{z}_{t}$ (lagged values and covariates), we produce one-step-ahead predictions $\hat{y}_{t+1}=f(\mathbf{z}_{t})$. An XAI method $\chi\in\{\text{SHAP},\text{LIME},\varnothing\}$ computes feature attributions $\boldsymbol{\phi}_{t+1}=\chi(f,\mathbf{z}_{t})$, where $\varnothing$ denotes no XAI (attribution-free baseline).
Tuples $\mathcal{R}=\{(\hat{y}_{i},\boldsymbol{\phi}_{i},E_{i},\mathbf{q}_{i})\}_{i=1}^{N}$ are generated in the pipeline, where $E_{i}=\text{LLM}(\mathcal{P};\hat{y}_{i},\boldsymbol{\phi}_{i},\mathbf{c})$ is the natural language explanation produced by prompting an LLM with strategy $\mathcal{P}$ and context $\mathbf{c}$ (model metrics, feature values), and $\mathbf{q}_{i}$ is the quality score from LLM-based evaluation. See Appendix C for a complete example of input context and generated explanation.
Our factorial design spans four factors (see Figure 2). Below we detail each factor and our evaluation framework; statistical analysis follows in Section 4.
3.2 Time-Series Forecasting Task
We use the UCI Individual Household Electric Power Consumption dataset (hebrail2012individual), containing 2,075,259 minute-level measurements of a French household’s electricity usage from December 2006 to November 2010. We chose this domain as a representative time-series forecasting task with connection to everyday decision-making: patterns are intuitive (e.g., higher consumption in winter due to heating) and LLMs can draw on common knowledge to produce meaningful explanations.
Data and Features.
We aggregate minute-level readings to weekly totals ($\approx$200 data points), balancing noise reduction with preserving temporal patterns. Features include lag values (past 7 weeks), ISO week number for seasonality, and holiday counts. We use a chronological 70/30 train/test split. See Appendix A for preprocessing details.
3.3 Forecasting Models
Model $\mathbf{R^{2}}$ MAE RMSE
XGBoost $0.69$ $20.55$ $25.03$
SARIMAX $0.55$ $20.12$ $29.91$
Random Forest $0.37$ $28.74$ $35.44$
MLP $0.21$ $33.29$ $39.69$
Table 1: Test-set performance metrics (sorted by R2).
We train three ML models (XGB, RF, MLP) that deliberately vary in prediction quality (Table 1), allowing us to test whether model performance affects NLE quality. We also include SARIMAX as a classical time-series counterpart with inherent interpretability through coefficients (no post-hoc XAI needed).
All models produce one-step-ahead forecasts under identical conditions: each prediction uses true observed values from all preceding weeks, not recursive model outputs. For ML models, lag features are pre-computed from actual consumption; for SARIMAX, we use a rolling approach that appends true values after each forecast. This ensures fair comparison across model types. See Appendix A for hyperparameters and visualizations.
Test Instance Selection.
To ensure representative coverage across the test period, we select three specific weeks using a tercile-based sampling strategy. We divide the 52-week test set into three equal segments (early, middle, late) and select one week from each tercile where absolute prediction errors rank consistently: $|\text{XGB}|<|\text{SARIMAX}|<|\text{RF}|<|\text{MLP}|$. This criterion ensures temporal diversity while focusing on instances where model quality differences are clearly expressed, with prediction errors ranging from $\approx$4% (XGBoost) to $\approx$36% (MLP). See Appendix A for specific instance details. For each of the three weeks and each model, we generate explanations across all XAI conditions, LLMs, and prompting strategies.
3.4 XAI Methods
For each ML model, we compute explanations using three conditions: SHAP (TreeExplainer for tree-based models, KernelExplainer for MLP), LIME and None (only prediction context without feature attributions). For SARIMAX, we use only the “none” condition since it provides inherent interpretability through coefficients. See Appendix A for implementation details.
3.5 Prompting Strategies
We select eight prominent strategies from the literature (sahoo2024systematic), spanning simple to complex approaches: zero-shot and few-shot (direct instruction with and without examples), chain-of-thought (zero- and few-shot variants, excluding DeepSeek-R1 which natively uses CoT), meta-prompting (structured 6-step template), reflexion (iterative self-critique), role-based (expert persona), and self-consistency (ensemble of multiple outputs). These range from single-pass methods to multi-iteration approaches. All prompts share a common structure: a system message establishing the task (help a non-technical user understand the forecast), and a human message providing context including model performance metrics, the predicted value, instance features, and XAI attributions when applicable. See Appendix B for complete templates.
We constrain explanations to $\leq$200 words and $\leq$6 bullet points, reflecting realistic user attention spans based on cognitive load research (mark2023attention; brysbaert2019reading) and controlling for verbosity bias (Section 2.4).
3.6 LLMs for NLE Generation
We use three LLMs spanning different scales and architectures: Llama-3-8B (grattafiori2024llama), a small open-source model run locally; GPT-4o (openai2024gpt4o), a large commercial model; and DeepSeek-R1 (guo2025deepseekr1), a 671B-parameter Mixture-of-Experts model with 37B active parameters, optimized for reasoning via reinforcement learning. This selection examines how model scale and reasoning specialization affect explanation quality. We use temperature $\tau=1.0$ for all strategies to encourage diverse outputs, except self-consistency’s synthesis step ($\tau=0.2$) which requires deterministic aggregation of multiple drafts into a single cohesive explanation (wang2023selfconsistency).
3.7 Evaluation Framework
We employ G-Eval (liu2023geval) for pointwise scoring with two judge models: GPT-4 and DeepSeek-R1. We exclude Llama as a judge due to its smaller parameter count – research shows that judge capability scales with model size, and weaker models produce less reliable evaluations (gu2024llmasjudge). Each dimension uses chain-of-thought evaluation steps followed by probability-weighted scoring (see Section 2.4). We define four evaluation dimensions grounded in XAI evaluation literature (hoffman2018metrics; miller2019explanation):
  • Accuracy: Faithfulness to provided XAI values – correct feature signs and approximate magnitudes.
  • Lay User Relevancy: Understandability for non-technical users. miller2019explanation notes that “practitioner satisfaction is fundamental to effective explanatory strategies.”
  • Expert Relevancy: Technical soundness for ML experts. ribeiro2016why assert that “good explanations should provide additional insights” enabling model assessment.
  • Explanation Helpfulness: Overall pragmatic usefulness. miller2019explanation notes individuals evaluate explanations “according to pragmatic usefulness.”
Each rubric uses a 1–5 scale with detailed criteria (see Appendix D). Following best practices from gu2024llmasjudge, scores are averaged across both judges to mitigate self-preference bias.
3.8 Corpus Summary
The factorial design yields the following corpus:
  • Main study: 3 test instances $\times$ 4 models $\times$ 3 XAI $\times$ 3 LLMs $\times$ 8 strategies
  • Exclusions: CoT strategies for DeepSeek-R1 (native reasoning); SHAP/LIME for SARIMAX (classical model)
  • Main corpus: 660 explanations
Table 2 summarizes generation efficiency (tokens and time) across all experimental factors.
Factor Level N Tokens Time (s)
[…]
Table 2: Generation efficiency by experimental factor (mean $\pm$ SD), ordered by tokens. Bold = best per factor. N varies by design: CoT excludes DeepSeek-R1; SARIMAX uses only “none” XAI.
4 Experiments and Results
For each research question, we test overall effects using one-way ANOVA. Beyond $p$-values, we report omega-squared ($\omega^{2}$) (kroes2023demystifying) as a measure of effect size – the proportion of variance in scores attributable to the factor (e.g., $\omega^{2}=.06$ for XAI means 6% of score variation is due to XAI choice). Following field2013discovering, we interpret: <.01<.01 negligible, .01–.06 small, .06–.14 medium, >.14>.14 large. Where ANOVA is significant, we conduct pairwise $t$-tests reporting Cohen’s $d$ (cohen1988statistical) – the standardized difference between group means. We interpret: <.2<.2 negligible, .2–.5 small, .5–.8 medium, >.8>.8 large.
Because we test five evaluation dimensions simultaneously, we apply Benjamini-Hochberg FDR correction (benjamini1995controlling) to control false discovery rates, reporting both raw $p$ and corrected $p_{\text{FDR}}$ values. We also fit factorial models (maxwell2017designing) controlling for other experimental factors (e.g., testing XAI effect while controlling for LLM and Strategy), and test for interactions to check whether effects vary across conditions.
Below we present the key findings; full statistical tables, robustness checks (Welch ANOVA, Games-Howell tests), and interaction analyses are in Appendix E.
4.1 RQ1: Does XAI Improve NLE Quality?
XAI Acc. Lay Exp. Help. Avg.
LIME 4.57 4.31 3.94 4.31 4.28
SHAP 4.53 4.32 3.97 4.30 4.28
None 4.49 4.31 3.78 4.38 4.24
Table 3: G-Eval scores by XAI condition, sorted by average (SARIMAX excluded). Acc./Lay/Exp./Help. are the four evaluation dimensions from §3.7; values are means across all samples per condition; Avg. is the mean of the four dimensions. Same format for all results tables. Bold = best per column.
[…]
Finding 1: XAI improves only Expert Relevancy (small effect, $d\approx 0.3$); other dimensions show no benefit. SHAP and LIME perform equivalently.
4.2 RQ2: Which LLM Produces the Best NLEs?
LLM Acc. Lay Exp. Help. Avg.
DeepSeek 4.71 4.72 4.45 4.78 4.66
GPT-4o 4.73 4.20 3.95 4.28 4.29
Llama-3 4.18 4.05 3.39 3.99 3.90
Table 4: G-Eval scores by generator LLM, sorted by average. Bold = best per column.
[…]
4.3 RQ3: Does ML Model Quality Affect NLEs?
Model Acc. Lay Exp. Help. Avg.
XGBoost 4.56 4.53 4.00 4.43 4.38
RF 4.53 4.29 3.82 4.27 4.23
MLP 4.50 4.12 3.86 4.29 4.19
Table 5: G-Eval scores by ML model, sorted by average. XGBoost (highest R2) yields best NLEs. Bold = best per column.
[…]
Finding 3: Better-performing ML models yield better NLEs (medium effect on Lay Relevancy, small on Expert/Helpfulness). Accuracy is unaffected by model quality.
4.4 RQ4: Does Inherent Interpretability Help?
Model Acc. Lay Exp. Help. Avg.
XGBoost 4.51 4.61 3.89 4.54 4.39
RF 4.46 4.26 3.68 4.29 4.17
MLP 4.50 4.07 3.76 4.32 4.16
SARIMAX 4.47 4.04 3.76 4.12 4.10
Table 6: NLE quality at XAI=“none” ($N=66$ each), sorted by average. Bold = best per column.
SARIMAX offers inherently interpretable coefficients (AR, MA, seasonal terms), unlike black-box ML models. Table 6 compares models at XAI=“none” (fair since SARIMAX cannot use SHAP/LIME).
SARIMAX yields the lowest scores despite having the second-highest R2. Significant Model effects emerge on Lay Relevancy ($\omega^{2}=.17$, medium) and Helpfulness ($\omega^{2}=.08$, small); both survive FDR. SARIMAX significantly underperforms XGBoost with large effects (d=0.98d=-0.98 to 1.16-1.16) and even underperforms lower-R2 models RF and MLP (d=0.37d=-0.37 to 0.44-0.44).
Finding 4: In our setting, SARIMAX yielded lower NLE quality than ML models despite higher prediction accuracy.
4.5 RQ5: Prompting Strategy Comparison
Strategy Acc. Lay Exp. Help. Avg.
Self-cons. 4.72 4.57 4.05 4.43 4.44
Zero-shot 4.70 4.36 4.00 4.36 4.35
Reflexion 4.74 4.39 3.96 4.31 4.35
Role-based 4.58 4.48 3.84 4.36 4.32
Few-shot 4.33 4.20 3.94 4.53 4.25
CoT Zero 4.71 4.19 3.64 4.12 4.17
Meta 4.22 3.97 3.80 4.06 4.01
CoT Few 4.13 4.02 3.68 4.22 4.01
Table 7: G-Eval scores by prompting strategy, sorted by average.
[…]
5 Discussion
5.1 What the Evaluation Dimensions Reveal
The four G-Eval dimensions capture distinct aspects of NLE quality. Our results show they respond differently to experimental factors, but LLM capability dominates all of them.
Accuracy measures factual correctness in representing provided information (numerical values, feature names, metrics). It is primarily determined by LLM choice ($\omega^{2}=.20$) and prompting strategy ($\omega^{2}=.14$). XAI condition has no effect: LLMs accurately report whatever information they receive, whether that includes feature attributions or not.
Lay Relevancy captures whether non-experts can understand the prediction rationale, assess trustworthiness, and derive actionable insights. It shows no XAI benefit but responds to ML model quality ($\omega^{2}=.12$) and prompting strategy ($\omega^{2}=.12$). The absent XAI effect aligns with research showing that lay users struggle to interpret feature attributions (salih2025perspective; alufaisan2020explainable). Better-performing models provide richer context for explanations; how the LLM frames that context matters equally.
[…]
Helpfulness evaluates whether users can judge prediction accuracy from cues like uncertainty ranges or error context. It shows the second-largest LLM effect ($\omega^{2}=.38$) with no XAI benefit. Communicating prediction reliability depends on LLM capability, not on whether XAI outputs are included.
5.2 SHAP vs. LIME
SHAP and LIME show no significant differences across any dimension. While these methods differ algorithmically (SHAP uses Shapley values from game theory, LIME fits local surrogate models), both produce similar feature attribution outputs for tabular data (salih2025perspective). The bottleneck for NLE quality appears to be the LLM’s translation of attributions into narrative, not the XAI method itself.
[…]
5.3 The Interpretability Paradox
SARIMAX yields lower NLE quality than ML models despite higher prediction accuracy (R2=.55 vs. .21–.37 for MLP/RF) and being inherently interpretable through its statistical coefficients. This is paradoxical: econometricians have interpreted classical time-series models for decades, and the very existence of XAI methods reflects that ML models are not inherently interpretable to humans. Yet LLMs produce better explanations for the opaque models than for the transparent ones.
We can only hypothesize about the cause, and it would need to be tested. One possibility is that LLM training data plays a role: feature-importance language (“temaperature contributed +0.3”) may be more common than statistical coefficient notation (AR, MA, seasonal terms). It may also be easier to translate XAI outputs into natural language, while AR, MA, and seasonal coefficients may be harder to explain accessibly, especially for lay users. If this is indeed an LLM limitation rather than a property of model types, fine-tuning on statistical content or improved prompting may close this gap.
5.4 Strategy Selection and Cost-Effectiveness
[…]
Chain-of-thought’s underperformance is surprising given its success on other tasks. One possibility is that explicit reasoning chains add verbosity without improving the final explanation, or that CoT is better suited to tasks requiring multi-step logical inference. Regardless of the mechanism, our results suggest CoT benefits are task-dependent, not universal.
6 Conclusion
We presented a systematic factorial study of factors affecting LLM-generated natural language explanations from XAI outputs. Across 4 ML models, 3 XAI conditions, 3 LLMs, and 8 prompting strategies, we generated and evaluated 660 explanations using G-Eval with dual LLM judges.
[…]
These findings challenge the assumption that sophisticated XAI methods automatically yield better explanations. The LLM’s capability to translate technical outputs matters more than the XAI method producing them. We offer testable hypotheses for the interpretability paradox.
Limitations
Scope. Our experiments focus on a single domain (household energy forecasting) using data from one household. Energy consumption is relatively intuitive – features like temperature and time-of-week have common-sense interpretations. Domains requiring specialized knowledge (e.g., genomics, materials science) or different data modalities (e.g., images, text) may yield different patterns.
Methods. We test only two XAI methods (SHAP and LIME), three LLMs (GPT-4o, DeepSeek-R1, Llama-3), and one classical model (SARIMAX). Other XAI approaches (e.g., attention-based, counterfactual, concept-based), other LLMs (e.g., Claude, Gemini), and other classical models (e.g., VAR, exponential smoothing) may behave differently. The interpretability paradox warrants replication with additional classical models. Furthermore, we use only prompt-based generation; fine-tuning or retrieval-augmented approaches may yield different results. Results may also be sensitive to prompt formulation.
Evaluation. We rely on LLM-based evaluation (G-Eval) rather than human annotators. Despite our bias mitigation strategies, LLM judges may have systematic biases not captured by our design. More importantly, we measure explanation quality but not utility. Research shows that explanation quality and user performance are often weakly correlated (schemmer2022should): users may over-rely on confident-sounding explanations or ignore technically accurate but complex ones. DeepSeek-R1 produces higher-quality explanations than GPT-4o, but whether users make better decisions with them remains untested. Similarly, the interpretability paradox concerns NLE quality, not whether users actually understand SARIMAX better when explained by humans versus LLMs.
Generalizability. Our findings may not generalize beyond time-series forecasting, as other domains (e.g., image classification, NLP) involve different XAI methods and explanation structures. The rapid pace of LLM development also means our specific rankings may not persist.
Appendix A Experimental Details
This section provides implementation details for data preprocessing, model training, and XAI computation.
A.1 Data Preprocessing
We use the UCI Individual Household Electric Power Consumption dataset (hebrail2012individual), containing 2,075,259 minute-level measurements from a household in Sceaux, France (near Paris), recorded from December 2006 to November 2010. The preprocessing pipeline: (1) handles missing values ($\approx$1.25%) via linear interpolation, (2) converts kilowatt-minutes to kilowatt-hours, and (3) aggregates to weekly totals anchored on Mondays (ISO week alignment):
Weekly Aggregation def load_weekly(file_path, country="FR"): df = pd.read_csv(file_path, sep=";", na_values="?") df["DateTime"] = pd.to_datetime(df["Date"] + " " + df["Time"]) # Handle missing values + convert kW-min -> kWh df["Global_active_power"] = ( df["Global_active_power"].astype(float).interpolate() * (1/60) ) # Weekly aggregation anchored on Monday (ISO alignment) w = df["Global_active_power"].resample("W-MON").sum() w = w.to_frame("True_Value").reset_index() # ISO week number (1-52) w["weekofyear"] = w["DateTime"].dt.isocalendar().week # Count holidays in each week fr_holidays = holidays.CountryHoliday(country) w["holiday_week_count"] = w["DateTime"].apply( lambda d: sum((d - pd.Timedelta(days=i)).date() in fr_holidays for i in range(7)) ) return w
Calendar features: weekofyear captures the ISO week number (1–52) for annual seasonality patterns (e.g., higher winter consumption). holiday_week_count counts French public holidays within each week using the holidays Python package—for each week ending on Monday, we count holidays in the preceding 7-day window.
Lag features: Seven autoregressive features (lag_1 through lag_7) capture the previous seven weeks of consumption:
[…]
A.2 Model Configurations
XGBoost.
Hyperparameters selected via 200-iteration random search with 4-fold time-series cross-validation:
XGBoost Configuration XGBRegressor( n_estimators = 1190, learning_rate = 0.00497, max_depth = 5, subsample = 0.805, colsample_bytree = 0.725, min_child_weight = 6, gamma = 0.248, reg_lambda = 0.530, early_stopping_rounds = 200 )
SARIMAX.
Orders selected via stepwise AICc search using pmdarima.auto_arima with: non-seasonal AR/MA orders $p,q\in[0,3]$, seasonal AR/MA orders $P,Q\in[0,2]$, forced seasonal differencing $D=1$, automatic non-seasonal differencing, Canova–Hansen seasonality test, and annual seasonality $m=52$. The selected configuration:
SARIMAX Configuration SARIMAX( endog = y_train, exog = train[["holiday_week_count"]], order = (2, 0, 1), seasonal_order = (1, 1, 0, 52), enforce_stationarity = False, enforce_invertibility = False, trend = ’c’ )
For test-set evaluation, we use rolling one-step-ahead forecasts: after each prediction, the true observed value is appended to the model state (without refitting) before forecasting the next week.
RF & MLP.
Deliberately weak configurations to span model quality:
Weak Model Configurations # Random Forest: shallow stumps RandomForestRegressor( n_estimators=10, max_depth=1, min_samples_leaf=25, max_features=0.08 ) # MLP: minimal capacity make_pipeline( StandardScaler(), MLPRegressor( hidden_layer_sizes=(8,), activation="tanh", solver="sgd", learning_rate_init=0.30, max_iter=30, n_iter_no_change=2 ) )
A.3 XAI Computation
SHAP.
TreeExplainer for tree-based models; KernelExplainer with 100-sample background for MLP:
SHAP Computation # Tree models (XGBoost, RandomForest) explainer = shap.TreeExplainer(model) shap_values = explainer.shap_values(X_test) base_value = explainer.expected_value # MLP (KernelExplainer) background = shap.sample(X_train, 100, random_state=42) explainer = shap.KernelExplainer(model.predict, background) shap_values = explainer.shap_values(X_test)
LIME.
Local surrogate with standardized features:
LIME Configuration explainer = LimeTabularExplainer( training_data = scaler.transform(X_train), feature_names = FEATURES, mode = "regression", discretize_continuous = False, sample_around_instance = True, kernel_width = 3, random_state = 42 ) # 8000 samples per explanation exp = explainer.explain_instance(row, predict_fn, num_samples=8000)
A.4 Prediction Visualization
Figure 3 shows test-set predictions for all four models against actual weekly consumption. XGBoost closely tracks the true series, while MLP exhibits high variance and poor trend-following.
Figure 3: Test-set predictions vs. actual weekly household energy consumption (kWh). XGBoost (R2=0.69) captures temporal patterns effectively; Random Forest (R2=0.37) shows moderate tracking; SARIMAX (R2=0.55) handles seasonality but misses peaks; MLP (R2=0.21) exhibits high variance with poor trend-following.
A.5 Test Instance Details
We select three test weeks using tercile-based sampling from the 52-week test set. The selection criterion requires absolute prediction errors to rank as: $|\text{XGB}|<|\text{SARIMAX}|<|\text{RF}|<|\text{MLP}|$, ensuring each instance demonstrates the quality gradient across models.
Tercile Date Week True (kWh)
Early 2009-11-09 46 208.20
Middle 2010-04-26 17 169.98
Late 2010-07-26 30 119.95
Table 8: Selected test instances.
Date Model Pred. Err. (%)
2009-11-09 XGBoost 175.4 -15.7
SARIMAX 249.3 +19.8
RF 164.1 -21.2
MLP 262.3 +26.0
2010-04-26 XGBoost 176.1 +3.6
SARIMAX 158.8 -6.6
RF 189.5 +11.5
MLP 231.8 +36.4
2010-07-26 XGBoost 125.0 +4.2
SARIMAX 97.3 -18.9
RF 156.3 +30.3
MLP 81.8 -31.8
Table 9: Predictions and errors per model.
Feature Nov 9 Apr 26 Jul 26
lag_1 184.7 189.2 127.8
lag_2 216.4 175.6 126.6
lag_3 197.2 196.4 150.3
lag_4 191.4 207.0 145.9
lag_5 166.3 184.6 168.4
lag_6 162.3 215.4 174.2
lag_7 189.8 137.4 163.9
weekofyear 46 17 30
holiday_count 0 0 0
Table 10: Feature values (lags in kWh).
Appendix B Prompt Templates
All prompting strategies share a common two-message structure: a System Message establishing the task and output constraints, and a Human Message providing the specific forecasting context.
B.1 Prompt Structure
System Message (Common) Interpret the time-series forecasting context that follows.
Your goal is to help a non-technical lay user. The user does not have a background in statistical models, machine learning, times-series or explainability methods (e.g. SHAP, LIME). The user needs to understand:
 -- why the model produced this forecast, and
 -- how much confidence can be placed in it.
OUTPUT RULES
 -- Write up to six bullet points -- no more.
 -- Keep the whole response <= 200 words.
 -- Use plain language, but key time-series terms such as lag, trend, seasonality, baseline, error are fine.
 -- Do not reveal code.
Human Message (Common Structure) The following is about time series data with a single-step ahead prediction, where the model predicts the next value in the time series based on previous observations.
Data Domain: Energy Consumption
Dataset Description:
 -- The dataset contains 2,075,259 measurements from a house in Sceaux, France (near Paris), Dec 2006--Nov 2010.
 -- Recorded at one-minute resolution, then resampled weekly.
 -- Lag_1..Lag_7, ISO week number and number of public holidays per week were added as features.
 -- Target: weekly global active power (kWh).
Model Used: [XGBoost / RandomForest / MLP / SARIMAX]
Model Performance:
 -- MAE: [value]
 -- RMSE: [value]
 -- R2: [value]
Prediction: [predicted value in kWh]
Instance Features or Context:
 lag_1: [value], lag_2: [value], ..., lag_7: [value]
 weekofyear: [1--52], holiday_week_count: [0--N]
[SHAP/LIME] values: (sorted by ||magnitude||, omitted for XAI=none)
 [feature]: [+/-value]
 The expected/base value for [SHAP/LIME]: [baseline]
(For LIME only:) Each LIME value is already $\beta_{i}\times x_{i}$: the additive kWh contribution of that feature for this instance.
SARIMAX-specific addition: For SARIMAX prompts, the human message additionally appends the full statsmodels model summary (coefficients, standard errors, diagnostics) followed by a note:
[Note] The SARIMAX order/seasonal_order were chosen via an auto_arima stepwise AICc search. Lagged demand effects are captured implicitly by the AR and MA terms; only ‘holiday_week_count’ enters as an explicit regressor.
B.2 Strategy-Specific Additions
Zero-shot.
Uses only the system prompt above with no additional instructions.
Few-shot.
Prepends four complete Human/Assistant example pairs demonstrating different domains and XAI methods. Below is one example; the remaining three are available in our GitHub repository.
Human Data Domain: Retail Sales
Dataset Description:
 -- 469 weekly sales totals from a mid-size supermarket chain in Madrid, Jan 2015 -- Dec 2023
 -- Originally daily cash-register data, aggregated to weekly sums
 -- Features: Lag_1...Lag_8, ISO week number, holidays per week, Promotion_Intensity (0--1)
 -- Target: weekly gross sales (thousand EUR)
Model Used: XGBRegressor
Model Performance: MAE: 12.3, RMSE: 16.8, R2: 0.874
Prediction: 242.70
Instance Features: lag_1: 238.4, lag_2: 245.1, ..., weekofyear: 48, holiday_week_count: 1, promotion_intensity: 0.65
SHAP values: promotion_intensity: +6.9, lag_1: +4.2, holiday_week_count: +3.5, lag_2: -2.8, lag_3: -1.1, ...
Base value: 235.1
Assistant -- SHAP starts from a baseline of ~235k EUR and shows how each feature nudges the forecast up (positive) or down (negative).
-- Promotion intensity (0.65) is the main driver, adding ~7k EUR.
-- Last-week sales (lag_1 = 238k EUR) add ~4k EUR, showing continued momentum.
-- One public holiday lifts demand by ~3.5k EUR, but the boost is modest.
-- Week 48 sits before Christmas; the tiny +0.4k EUR seasonal bump hints demand hasn’t peaked yet.
-- Negative pulls from lags 2-4 temper the forecast; 242.7k EUR is plausible within the +/-16.8k EUR RMSE.
CoT Zero-shot.
Appends to system prompt:
Think step-by-step.
CoT Few-shot.
Combines few-shot examples with CoT instruction.
Role-based.
Appends to system prompt:
You are a seasoned data-scientist and approachable teacher.
You love turning complex numbers into everyday language so
home-owners can make smart, confident decisions.
Meta-prompting.
Uses a structured 6-step reasoning template prepended to the human message:
Let’s think step by step:
 -- Context: Summarise the problem domain and briefly name the XAI method.
 -- Data Insights: Note dataset size, key features, possible biases.
 -- Model/Parameters: Mention any parameter that affects performance.
 -- Performance Metrics: Explain MAE, RMSE, R2.
 -- XAI / Feature Importance: (1) Top factors. (2) Why they matter. (3) Surprises.
 -- Summary: Tie the insights into a final answer.
Reflexion.
Iterative self-improvement loop with up to 6 trials. Each iteration: (1) generate draft, (2) grade via GPT-4o, (3) self-reflect. Stops early if all scores $\geq$4. Memory retains last 3 reflections.
Grading Prompt (GPT-4o) You are evaluating a time-series forecast explanation for a lay reader.
Ground-truth facts
Prediction: [value]
Top features $\pm$ values: [XAI attributions]
Explanation to grade
[draft explanation] >>>>>>
Rate each criterion 1--5 (briefly justify $\leq$15 words):
 1. Accuracy / Faithfulness
 2. Coverage of key drivers
 3. Confidence communication
 4. Clarity for a lay reader
Respond with one valid JSON line:
{"accuracy": x, "coverage": y, "confidence": z, "clarity": w, "justifications": {...}}
Reflection Prompt Outcome: [SUCCESS/FAIL].
Give ONE short sentence advising yourself how to improve accuracy, coverage, confidence, clarity next time.
Self-consistency.
Generates 7 independent drafts at temperature $\tau$=1.0, then synthesizes into one cohesive explanation at $\tau$=0.2.
Synthesis Prompt ($\tau$=0.2) Below are multiple candidate explanations produced by the model for the same input. Please create a single, cohesive explanation that:
 -- Retains the most insightful sentences from each.
 -- Resolves contradictions.
 -- Still obeys the format instructions.
=== CANDIDATES ===
--- Candidate #1 ---
[draft 1]
--- Candidate #2 ---
[draft 2]
... (7 total)
Appendix C Example Input and Output
Below is a complete example from our energy consumption domain using zero-shot prompting with SHAP.
C.1 Input Context
Input (Zero-shot with SHAP) Data Domain: Household Energy Consumption
Dataset Description:
 -- 200 weekly power consumption totals from a French household, Dec 2006 -- Nov 2010
 -- Originally minute-level data, aggregated to weekly sums (kWh)
 -- Features: lag_1 through lag_7 (past 7 weeks), weekofyear (1--52), holiday_count
 -- Target: weekly global active power (kWh)
Model Used: XGBoost Regressor
Model Performance: MAE: 20.55, RMSE: 25.03, R2: 0.69
Prediction: 175.44 kWh
Instance Features:
 lag_1: 163.10, lag_2: 158.90, lag_3: 171.20, ..., weekofyear: 42, holiday_count: 1
SHAP Values:
 lag_1: +12.34, lag_2: -5.21, lag_3: +3.45, weekofyear: -8.12, holiday_count: +2.89
Base Value: 152.00 kWh
C.2 Generated Explanation
Generated NLE (Zero-shot) -- Predicted consumption: The model forecasts 175.44 kWh for next week, about 15% above the typical weekly average of 152 kWh.
-- Main driver: Last week’s consumption (163.10 kWh) is the strongest factor pushing the forecast up by about 12 kWh. When recent usage is high, the model expects that pattern to continue.
-- Seasonal adjustment: Week 42 (mid-October) typically has lower consumption, which pulls the forecast down by about 8 kWh. Autumn weather may reduce heating/cooling needs.
-- Holiday effect: One public holiday this week adds roughly 3 kWh to the forecast, as holidays often correlate with more time spent at home.
-- Model reliability: With an R2 of 0.69, the model explains about 69% of the variation in weekly consumption. The average error is around 20--25 kWh, so actual consumption could reasonably fall between 150--200 kWh.
-- Bottom line: Expect moderately high consumption driven by recent usage patterns, partially offset by the season. Plan for some uncertainty given the model’s typical error range.
Appendix D Evaluation Rubrics
Each G-Eval dimension uses chain-of-thought evaluation steps followed by a 1–5 score.
D.1 Accuracy Rubric
Accuracy (1–5) You are evaluating the ACCURACY of a time-series forecast explanation.
Context: Time series data with single-step ahead prediction of weekly global active power in kWh.
Scoring:
 5: All provided information is accurately represented (numerical values, feature names, relationships, metrics)
 4: Minor inaccuracies in non-critical details but core provided facts are correct
 3: Some factual errors in representing provided data but main message intact
 2: Multiple factual errors that misrepresent the provided information
 1: Major hallucinations or significant misrepresentation of provided data
Chain of Thought Steps:
 1. Identify what information was actually provided in the context
 2. Check if mentioned numerical values match the provided data
 3. Verify that any referenced feature names and values are correctly stated
 4. Confirm that any mentioned model performance metrics are accurately reported
 5. Assess if relationships described are factually correct based on provided information
 6. Rate overall factual accuracy of representing the PROVIDED information only
Important: Do not penalize explanations for missing information if none was provided in the context. Only evaluate accuracy of what information was actually given.
D.2 Lay User Relevancy Rubric
Lay User Relevancy (1–5) You are a NON-TECHNICAL USER evaluating a forecast explanation.
You have NO background in machine learning, statistics, or technical methods. You just want to understand why the system made this prediction and whether you should trust it.
Scoring:
 5: Extremely relevant and helpful for understanding and trusting the prediction
 4: Very relevant with useful information for my needs
 3: Somewhat relevant but could be more helpful
 2: Limited relevance to what I need to know
 1: Not relevant or confusing for a non-technical user
Chain of Thought Steps:
 1. Does this explanation help me understand WHY the prediction was made?
 2. Can I determine if I should trust this prediction based on the explanation?
 3. Is the information presented in a way I can understand?
 4. Does this give me actionable insights for my decision-making?
 5. Rate overall relevancy for a lay user
D.3 Expert Relevancy Rubric
Expert Relevancy (1–5) You are a MACHINE LEARNING EXPERT evaluating a forecast explanation.
You have deep knowledge of machine learning models and time-series forecasting methods.
Scoring:
 5: Extremely relevant with appropriate technical depth for the information provided
 4: Very relevant with good technical information
 3: Somewhat relevant but missing some key technical aspects
 2: Limited technical relevance
 1: Not technically relevant or misleading
Chain of Thought Steps:
 1. Does the explanation provide sufficient technical reasoning given the available information?
 2. Can I assess model behavior and prediction reliability from this explanation?
 3. Is the technical content accurate and well-presented?
 4. Does this meet expert needs given available information?
 5. Rate overall technical relevancy
D.4 Explanation Helpfulness Rubric
Explanation Helpfulness (1–5) You are a NON-TECHNICAL USER evaluating how useful a forecast explanation is for judging prediction accuracy.
You have NO background in machine learning or statistics. You just want to know if the explanation gives you enough information to decide how close the prediction is to the true value.
Scoring:
 5: Makes it very clear how accurate the prediction is
 4: Gives good guidance about prediction accuracy
 3: Provides some helpful information
 2: Gives limited guidance
 1: Does not help assess accuracy at all
Chain of Thought Steps:
 1. Check clarity: Is the explanation understandable for a lay user (plain language, concrete references)?
 2. Check evidence: Does it include cues that inform accuracy (e.g., recent error rates, uncertainty ranges, comparison to history/seasonality, drivers affecting reliability)?
 3. Check actionability: Can a lay user form a reasonable judgment about closeness from the information given?
 4. Penalize jargon-only or vague statements that do not improve judgment
 5. Assign a score using the 1--5 scale above
Appendix E Statistical Analysis Details
This appendix provides complete statistical analysis for each research question. To ensure rigor and reproducibility, we apply a standardized analysis protocol across all RQs (except RQ4, which requires a specialized design due to SARIMAX constraints).
Evaluation Dimensions.
We evaluate NLE quality across four dimensions: Accuracy (faithfulness to XAI outputs), Lay Relevancy (accessibility for non-experts), Expert Relevancy (technical depth), and Helpfulness (overall usefulness). Each dimension uses a 1–5 scale; scores are averaged across both judges (GPT-4 and DeepSeek-R1) to mitigate self-preference bias.
Standardized Analysis Pipeline.
For each RQ, we apply the following sequence:
  1. 1.
    One-way ANOVA with omega-squared ($\omega^{2}$) effect sizes (kroes2023demystifying). We report $F$-statistics, raw $p$-values, FDR-corrected $p$-values (benjamini1995controlling), and effect size classifications.
  2. 2.
    Pairwise comparisons using independent $t$-tests with Cohen’s $d$ effect sizes (cohen1988statistical) and Benjamini-Hochberg FDR correction applied within each dimension.
  3. 3.
    Welch ANOVA as robustness check (welch1951comparison), which does not assume equal variances across groups – important given our unbalanced designs (e.g., CoT strategies exclude DeepSeek-R1).
  4. 4.
    Games-Howell post-hoc tests (games1976pairwise) for significant dimensions, which are robust to unequal variances and sample sizes.
  5. 5.
    Factorial linear model (Type II SS) controlling for other experimental factors, testing whether the main effect persists after accounting for confounds.
  6. 6.
    Interaction analysis testing whether the main factor’s effect varies across levels of other factors (e.g., does XAI effect differ by LLM?). We test all two-way interactions: Factor $\times$ LLM, Factor $\times$ Model, Factor $\times$ XAI, and Factor $\times$ Strategy (where applicable).
  7. 7.
    Simple effects analysis for significant interactions, examining the main factor’s effect within each level of the moderating factor.
Effect Size Interpretation.
Following cohen1988statistical and kroes2023demystifying:
  • $\omega^{2}$ (omnibus): <.01<.01 negligible, .01.01.06.06 small, .06.06.14.14 medium, >.14>.14 large
  • Cohen’s $d$ / Hedges’ $g$ (pairwise): <.2<.2 negligible, .2.2.5.5 small, .5.5.8.8 medium, >.8>.8 large
Multiple Comparison Correction.
We apply Benjamini-Hochberg FDR correction (benjamini1995controlling) to control false discovery rate at $\alpha=.05$. FDR correction is applied: (1) across dimensions within each ANOVA, and (2) within each dimension for pairwise comparisons.
RQ4 Exception.
RQ4 (SARIMAX vs. ML models) uses a specialized design because SARIMAX only supports XAI=“none.” Fair comparisons filter to this condition ($N=264$), and we focus on testing the “interpretability paradox” hypothesis rather than following the full interaction protocol.
RQ5 Supplementary Analysis.
For RQ5 (prompting strategy), we supplement pointwise G-Eval with listwise ranking evaluation. Both judges ranked all 8 strategies for 60 matched sets, validated via Friedman test and correlation with G-Eval scores.
E.1 RQ1: XAI Effect
Analysis excludes SARIMAX (only supports XAI=“none”), yielding $N=594$ with balanced groups (198 per XAI condition).
ANOVA Results.
Table 11 shows the omnibus test for XAI effect across all four dimensions.
Dimension F p pFDR $\omega^{2}$ Effect Sig.
Accuracy 1.07 .342 .456 .000 negl.
Lay Relevancy 0.00 .999 .999 .000 negl.
Expert Relevancy 6.32 .002 .008 .018 small **
Helpfulness 1.57 .209 .418 .002 negl.
Table 11: One-way ANOVA for XAI effect (RQ1). Only Expert Relevancy is significant.
Pairwise Comparisons.
Table 12 shows pairwise $t$-tests with Cohen’s $d$ effect sizes and FDR-corrected $p$-values for all dimensions.
Dimension Comparison $\Delta$ $d$ Effect p pFDR
Accuracy SHAP vs None +0.04 +0.08 negl. .445 .479
LIME vs None +0.08 +0.15 negl. .149 .448
SHAP vs LIME -0.04 -0.07 negl. .479 .479
Lay Rel. SHAP vs None +0.00 +0.01 negl. .960 .996
LIME vs None +0.00 +0.00 negl. .996 .996
SHAP vs LIME +0.00 +0.00 negl. .965 .996
Expert Rel. SHAP vs None +0.19 +0.33 small .001 .003*
LIME vs None +0.16 +0.28 small .006 .008*
SHAP vs LIME +0.03 +0.06 negl. .558 .558
Helpfulness SHAP vs None -0.08 -0.16 negl. .108 .221
LIME vs None -0.07 -0.15 negl. .147 .221
SHAP vs LIME -0.01 -0.02 negl. .822 .822
Table 12: Pairwise comparisons for RQ1 (XAI effect). FDR correction applied within each dimension.
Robustness Check: Welch ANOVA.
Table 13 shows Welch’s ANOVA, which is robust to heterogeneous variances. Results confirm standard ANOVA findings.
Dimension F p Sig.
Accuracy 1.04 .354
Lay Relevancy 0.00 .999
Expert Relevancy 6.45 .002 **
Helpfulness 1.60 .203
Table 13: Welch ANOVA for RQ1 (robust to unequal variances).
Post-hoc Tests.
Table 14 shows Games-Howell post-hoc results for Expert Relevancy (the only significant dimension).
A B Mean(A) Mean(B) Diff SE p Hedges’ $g$
LIME None 3.94 3.78 +0.16 0.057 .016 +0.28
LIME SHAP 3.94 3.97 -0.03 0.059 .827 -0.06
None SHAP 3.78 3.97 -0.19 0.058 .003 -0.33
Table 14: Games-Howell for Expert Relevancy (RQ1). Robust to unequal variances.
Factorial Model.
Table 15 shows XAI effect when controlling for LLM and Strategy (Type II SS).
Dimension XAI F XAI p Sig.
Accuracy 1.57 .209
Lay Relevancy 0.00 .998
Expert Relevancy 12.93 <<.001 ***
Helpfulness 2.98 .051
Table 15: Factorial model: XAI effect controlling for LLM and Strategy (RQ1).
Interaction Analysis.
Tables 1618 show all two-way interactions for the XAI factor.
Dimension F p $\omega^{2}$ Sig.
Accuracy 2.25 .062 .008
Lay Relevancy 3.05 .017 .014 *
Expert Relevancy 2.68 .031 .011 *
Helpfulness 2.00 .093 .007
Table 16: XAI $\times$ LLM interaction tests (RQ1).
Dimension F p $\omega^{2}$ Sig.
Accuracy 0.39 .818 .000
Lay Relevancy 0.97 .424 .000
Expert Relevancy 0.12 .974 .000
Helpfulness 0.91 .456 .000
Table 17: XAI $\times$ Model interaction tests (RQ1). No significant interactions.
Dimension F p $\omega^{2}$ Sig.
Accuracy 1.32 .189 .008
Lay Relevancy 0.67 .800 .000
Expert Relevancy 1.05 .403 .001
Helpfulness 0.40 .976 .000
Table 18: XAI $\times$ Strategy interaction tests (RQ1). No significant interactions.
Simple Effects.
Table 19 shows XAI effect within each LLM for dimensions with significant interactions.
Dimension LLM F p SHAP LIME None
Expert Rel. GPT-4o 16.01 <<.001* 4.13 4.01 3.79
Llama-3 1.36 .258 3.40 3.49 3.35
DeepSeek 4.64 .011* 4.51 4.44 4.33
Lay Rel. GPT-4o 0.11 .893 4.23 4.23 4.26
Llama-3 1.40 .248 4.01 4.03 4.15
DeepSeek 7.05 .001* 4.84 4.81 4.61
Table 19: Simple effects: XAI effect within each LLM (RQ1).
E.2 RQ2: LLM Effect
Analysis uses full dataset ($N=660$). Note unbalanced design: DeepSeek-R1 has 180 samples (CoT strategies excluded as native reasoning model), while GPT-4o and Llama-3 have 240 each.
ANOVA Results.
Table 20 shows the omnibus test for LLM effect. All four dimensions show highly significant effects with large effect sizes.
Dimension F p pFDR $\omega^{2}$ Effect Sig.
Accuracy 81.32 <<.001 <<.001 .196 large ***
Lay Relevancy 107.85 <<.001 <<.001 .245 large ***
Expert Relevancy 332.15 <<.001 <<.001 .501 large ***
Helpfulness 206.35 <<.001 <<.001 .384 large ***
Table 20: One-way ANOVA for LLM effect (RQ2). Large effects on all dimensions.
Pairwise Comparisons.
Table 21 shows pairwise $t$-tests with Cohen’s $d$ and FDR-corrected $p$-values.
Dimension Comparison $\Delta$ $d$ Effect p pFDR
Accuracy DeepSeek vs GPT -0.02 -0.04 negl. .680 .680
DeepSeek vs Llama +0.53 +0.91 large <<.001 <<.001*
GPT vs Llama +0.54 +0.95 large <<.001 <<.001*
Lay Rel. DeepSeek vs GPT +0.51 +1.17 large <<.001 <<.001*
DeepSeek vs Llama +0.67 +1.40 large <<.001 <<.001*
GPT vs Llama +0.16 +0.31 small .001 .001*
Expert Rel. DeepSeek vs GPT +0.50 +1.33 large <<.001 <<.001*
DeepSeek vs Llama +1.06 +2.50 large <<.001 <<.001*
GPT vs Llama +0.56 +1.24 large <<.001 <<.001*
Helpfulness DeepSeek vs GPT +0.50 +1.55 large <<.001 <<.001*
DeepSeek vs Llama +0.79 +1.91 large <<.001 <<.001*
GPT vs Llama +0.28 +0.66 medium <<.001 <<.001*
Table 21: Pairwise comparisons for RQ2 (LLM effect). FDR correction applied within each dimension.
Robustness Check: Welch ANOVA.
Table 22 shows Welch’s ANOVA (robust to unequal variances and sample sizes). Results confirm standard ANOVA.
Dimension F p Sig.
Accuracy 59.81 <<.001 ***
Lay Relevancy 128.21 <<.001 ***
Expert Relevancy 363.22 <<.001 ***
Helpfulness 261.40 <<.001 ***
Table 22: Welch ANOVA for RQ2 (robust to unequal variances/samples).
Post-hoc Tests.
Games-Howell for Expert Relevancy (strongest effect): DeepSeek vs GPT ($\Delta=0.50$, $p<.001$, Hedges’ $g=1.33$), DeepSeek vs Llama ($\Delta=1.06$, $p<.001$, $g=2.50$), GPT vs Llama ($\Delta=0.56$, $p<.001$, $g=1.24$).
Factorial Model.
Table 23 shows LLM effect when controlling for XAI and Strategy (Type II SS). Effects remain highly significant.
Dimension LLM F LLM p Sig.
Accuracy 95.70 <<.001 ***
Lay Relevancy 112.58 <<.001 ***
Expert Relevancy 334.24 <<.001 ***
Helpfulness 220.51 <<.001 ***
Table 23: Factorial model: LLM effect controlling for XAI and Strategy (RQ2).
Interaction Analysis.
Tables 2426 show all two-way interactions for the LLM factor.
Dimension F p $\omega^{2}$ Sig.
Accuracy 18.79 <<.001 .276 ***
Lay Relevancy 19.81 <<.001 .287 ***
Expert Relevancy 23.23 <<.001 .323 ***
Helpfulness 33.18 <<.001 .408 ***
Table 24: LLM $\times$ Strategy interaction tests (RQ2). All significant.
Dimension F p $\omega^{2}$ Sig.
Accuracy 1.60 .172 .004
Lay Relevancy 3.17 .014 .013 *
Expert Relevancy 3.63 .006 .016 **
Helpfulness 2.22 .066 .007
Table 25: LLM $\times$ XAI interaction tests (RQ2). Small effects on 2 dimensions.
Dimension F p $\omega^{2}$ Sig.
Accuracy 1.08 .374 .001
Lay Relevancy 1.06 .387 .001
Expert Relevancy 3.68 .001 .024 **
Helpfulness 0.38 .891 .000
Table 26: LLM $\times$ Model interaction tests (RQ2). Expert Relevancy shows significant interaction.
Simple Effects.
LLM effect within each Strategy for Helpfulness (representative dimension): All strategies show significant LLM effects ($p<.05$). DeepSeek-R1 consistently achieves highest scores when available (excluded from CoT strategies). Within CoT strategies (GPT vs Llama only), GPT-4o outperforms Llama-3.
E.3 RQ3: ML Model Effect
Analysis uses ML models only ($N=594$): XGBoost ($N=198$), Random Forest ($N=198$), MLP ($N=198$). SARIMAX is analyzed separately in RQ4.
ANOVA Results.
Table 27 shows the omnibus test for Model effect across three ML models.
Dimension F p pFDR $\omega^{2}$ Effect Sig.
Accuracy 0.60 .550 .550 .000 negl.
Lay Relevancy 40.12 <<.001 <<.001 .117 medium ***
Expert Relevancy 5.41 .005 .006 .015 small **
Helpfulness 5.70 .004 .005 .016 small **
Table 27: One-way ANOVA for ML Model effect (RQ3). $N=594$, $df=(2,591)$.
Pairwise Comparisons.
Table 28 shows pairwise $t$-tests with Cohen’s $d$ effect sizes and FDR-corrected $p$-values for all dimensions.
Dimension Comparison $\Delta$ $d$ Effect p pFDR
Lay Rel. XGB vs RF +0.25 +0.48 small <<.001 <<.001*
XGB vs MLP +0.41 +0.82 large <<.001 <<.001*
RF vs MLP +0.16 +0.30 small .003 .004*
Expert Rel. XGB vs RF +0.18 +0.33 small .001 .003*
XGB vs MLP +0.14 +0.24 small .024 .036*
RF vs MLP -0.04 -0.07 negl. .452 .452
Helpfulness XGB vs RF +0.16 +0.32 small .001 .004*
XGB vs MLP +0.14 +0.29 small .007 .014*
RF vs MLP -0.02 -0.04 negl. .681 .681
Table 28: Pairwise comparisons for RQ3 (Model effect). FDR correction applied within each dimension.
Robustness Check: Welch ANOVA.
Table 29 confirms results using Welch’s ANOVA (robust to variance heterogeneity).
Dimension F p Sig.
Accuracy 0.59 .554
Lay Relevancy 40.28 <<.001 ***
Expert Relevancy 5.28 .005 **
Helpfulness 5.68 .004 **
Table 29: Welch ANOVA for RQ3 (ML models only).
Games-Howell Post-hoc.
Robust to unequal variances. Significant pairs: Lay Relevancy (3): XGB>>RF ($g=0.48$), XGB>>MLP ($g=0.82$), RF>>MLP ($g=0.30$). Expert Relevancy (1): XGB>>RF ($g=0.33$). Helpfulness (2): XGB>>MLP ($g=0.29$), XGB>>RF ($g=0.32$). Confirms Tukey HSD findings.
Factorial Model.
Table 30 shows Model effect controlling for LLM, XAI, and Strategy (Type II SS).
Dimension F p $\eta^{2}_{p}$ Sig.
Accuracy 1.02 .360 .004
Lay Relevancy 51.94 <<.001 .152 ***
Expert Relevancy 11.65 <<.001 .039 ***
Helpfulness 11.89 <<.001 .039 ***
Table 30: Factorial model: Model effect controlling for confounds (RQ3).
Interaction Tests.
Tables 3133 show all two-way interactions for the Model factor.
Dimension F p $\eta^{2}_{p}$ Sig.
Accuracy 0.67 .615 .005
Lay Relevancy 1.12 .347 .008
Expert Relevancy 1.66 .157 .011
Helpfulness 0.16 .958 .001
Table 31: Model $\times$ LLM interaction tests (RQ3). No significant interactions.
Dimension F p $\eta^{2}_{p}$ Sig.
Accuracy 0.39 .818 .003
Lay Relevancy 0.97 .424 .007
Expert Relevancy 0.12 .974 .001
Helpfulness 0.91 .456 .006
Table 32: Model $\times$ XAI interaction tests (RQ3). No significant interactions.
Dimension F p $\eta^{2}_{p}$ Sig.
Accuracy 0.68 .794 .016
Lay Relevancy 0.77 .704 .019
Expert Relevancy 0.44 .960 .011
Helpfulness 0.50 .936 .012
Table 33: Model $\times$ Strategy interaction tests (RQ3). No significant interactions.
No significant interactions (all $p>.05$), indicating Model effects are consistent across LLMs and Strategies.
Simple Effects.
Model effect within each LLM for Lay Relevancy (representative dimension): GPT ($F=29.0$, $p<.001$, $\omega^{2}=.21$), DeepSeek ($F=15.0$, $p<.001$, $\omega^{2}=.15$), Llama ($F=8.7$, $p<.001$, $\omega^{2}=.07$). All LLMs show significant Model effects ($p<.05$) for Lay Relevancy and Expert Relevancy. For Helpfulness, GPT and DeepSeek show significant effects.
Model R2 vs NLE Quality.
Spearman $\rho=1.00$ between model R2 and overall NLE score ($n=3$, descriptive only). Higher-performing models yield better explanations.
E.4 RQ4: SARIMAX (Classical Model)
Analysis compares SARIMAX (classical time series, $N=66$) against ML models. SARIMAX only supports XAI=“none” in our design, so fair comparisons filter to XAI=“none” only ($N=264$).
ANOVA at XAI=“none”.
Table 34 shows Model effect when all models have XAI=“none”.
Dimension F p pFDR $\omega^{2}$ Effect Sig.
Accuracy 0.11 .952 .952 .000 negl.
Lay Relevancy 19.24 <<.001 <<.001 .172 large ***
Expert Relevancy 1.40 .244 .305 .004 negl.
Helpfulness 8.89 <<.001 <<.001 .082 medium ***
Table 34: ANOVA for Model effect at XAI=“none” (RQ4). $N=264$, $df=(3,260)$.
Post-hoc: SARIMAX vs ML Models.
Table 35 compares SARIMAX to each ML model at XAI=“none”.
Dimension vs $\Delta$ $d$ p Sig.
Lay Relevancy XGB -0.57 -1.16 <<.001 ***
Lay Relevancy RF -0.23 -0.44 .012 *
Lay Relevancy MLP -0.04 -0.07 .674
Helpfulness XGB -0.42 -0.98 <<.001 ***
Helpfulness RF -0.17 -0.37 .037 *
Helpfulness MLP -0.19 -0.41 .020 *
Table 35: SARIMAX vs each ML model at XAI=“none” (N=66 each). FDR-corrected.
Testing the R2 Hypothesis.
If prediction quality (R2) drove NLE quality, SARIMAX (R2=0.55) should outperform RF (R2=0.37) and MLP (R2=0.21). Table 36 tests this hypothesis.
[…]
SARIMAX vs MLP Helpfulness +0.34 -0.19 -0.41 .020*
SARIMAX vs RF Lay Relevancy +0.18 -0.23 -0.44 .012*
SARIMAX vs RF Helpfulness +0.18 -0.17 -0.37 .037*
[…]
Despite having higher R2, SARIMAX yields significantly worse NLEs on multiple dimensions. This falsifies the R2 hypothesis: prediction quality does not drive NLE quality.
SARIMAX vs Pooled ML.
Comparing SARIMAX ($N=66$) against all ML models ($N=594$, all XAI conditions): significant deficits on Lay Relevancy ($\Delta=-0.28$, d=0.51d=-0.51, $p_{\text{FDR}}<.001$) and Helpfulness ($\Delta=-0.21$, d=0.42d=-0.42, $p_{\text{FDR}}=.003$). Expert Relevancy shows marginal deficit (d=0.24d=-0.24, $p=.07$). Accuracy unaffected.
LLM Effect Within SARIMAX.
Strong LLM effects: Expert Relevancy ($F=71.4$, $\omega^{2}=.68$), Helpfulness ($F=18.0$, $\omega^{2}=.34$), Accuracy ($F=15.6$, $\omega^{2}=.31$), Lay Relevancy ($F=7.5$, $\omega^{2}=.17$). DeepSeek-R1 achieves highest overall score (4.09), followed by GPT-4o (3.72), then Llama-3 (3.41).
Strategy Effect Within SARIMAX.
Modest Strategy effects on Lay Relevancy ($F=3.4$, $p=.005$, $\omega^{2}=.20$) and Accuracy ($F=2.5$, $p=.028$, $\omega^{2}=.13$). Top strategies: self-consistency (3.88), reflexion (3.86), role-based (3.76).
The Interpretability Paradox.
SARIMAX has inherently interpretable coefficients (AR, MA, seasonal), yet yields worse NLEs than black-box ML models – even models with lower prediction quality. If R2 were the driver, expected ranking would be: XGBoost (0.69) >> SARIMAX (0.55) >> RF (0.37) >> MLP (0.21). Actual NLE ranking: XGBoost >> RF >> MLP >> SARIMAX. This reveals that model type matters more than prediction quality: LLMs struggle to translate statistical coefficients (AR, MA terms) into accessible narratives, regardless of how accurate the predictions are.
E.5 RQ5: Prompting Strategy Effect
Analysis uses full dataset ($N=660$). Strategies have unbalanced sample sizes due to design constraints: CoT strategies excluded DeepSeek-R1 (native reasoning model), yielding $N=60$ for CoT-zero and CoT-few vs. $N=90$ for other strategies.
ANOVA Results.
Table 37 shows the omnibus test for Strategy effect across all four dimensions.
Dimension F p pFDR $\omega^{2}$ Effect Sig.
Accuracy 16.75 <<.001 <<.001 .143 large ***
Lay Relevancy 14.28 <<.001 <<.001 .123 medium ***
Expert Relevancy 4.80 <<.001 <<.001 .039 small ***
Helpfulness 8.86 <<.001 <<.001 .077 medium ***
Table 37: One-way ANOVA for Strategy effect (RQ5). $N=660$, $df=(7,652)$.
Robustness Check: Welch ANOVA.
Table 38 shows Welch’s ANOVA, which is robust to heterogeneous variances and unequal sample sizes. Results confirm standard ANOVA findings.
Dimension F p Sig.
Accuracy 14.42 <<.001 ***
Lay Relevancy 16.47 <<.001 ***
Expert Relevancy 5.46 <<.001 ***
Helpfulness 9.79 <<.001 ***
Table 38: Welch ANOVA for RQ5 (robust to unequal variances/samples).
Games-Howell Post-hoc.
Table 39 shows significant pairwise comparisons (robust to unequal variances). Only pairs with $p_{\text{adj}}<.05$ shown.
Dimension Comparison $\Delta$ Hedges’ $g$ padj Sig.
Accuracy Self-cons. vs CoT-few +0.59 +1.10 <<.001 ***
Accuracy Reflexion vs CoT-few +0.61 +1.14 <<.001 ***
Accuracy Zero-shot vs CoT-few +0.57 +1.10 <<.001 ***
Accuracy CoT-zero vs CoT-few +0.58 +1.04 <<.001 ***
Accuracy Self-cons. vs Meta +0.50 +0.93 <<.001 ***
Lay Rel. Self-cons. vs Meta +0.60 +1.23 <<.001 ***
Lay Rel. Self-cons. vs CoT-few +0.55 +1.02 <<.001 ***
Lay Rel. Role-based vs Meta +0.51 +1.01 <<.001 ***
Lay Rel. Reflexion vs Meta +0.42 +0.83 .001 **
Expert Rel. Self-cons. vs CoT-few +0.51 +0.85 .003 **
Helpfulness Self-cons. vs Meta +0.37 +0.80 .005 **
Helpfulness Few-shot vs Meta +0.47 +0.97 <<.001 ***
Table 39: Games-Howell significant pairs for Strategy (RQ5). FDR-corrected.
Factorial Model.
Table 40 shows Strategy effect when controlling for LLM and XAI (Type II SS). Effects remain highly significant.
Dimension Strategy F p Sig.
Accuracy 20.80 <<.001 ***
Lay Relevancy 15.66 <<.001 ***
Expert Relevancy 3.55 .001 **
Helpfulness 12.06 <<.001 ***
Table 40: Factorial model: Strategy effect controlling for LLM and XAI (RQ5).
Two-way Interactions.
Tables 4143 test all two-way interactions with Strategy.
Dimension F p $\omega^{2}$ Sig.
Accuracy 18.79 <<.001 .276 ***
Lay Relevancy 19.81 <<.001 .287 ***
Expert Relevancy 23.23 <<.001 .323 ***
Helpfulness 33.18 <<.001 .408 ***
Table 41: Strategy $\times$ LLM interaction (RQ5). All dimensions show significant large effects.
Dimension F p $\omega^{2}$ Sig.
Accuracy 1.25 .231 .005
Lay Relevancy 0.79 .680 .000
Expert Relevancy 0.93 .522 .000
Helpfulness 0.35 .987 .000
Table 42: Strategy $\times$ XAI interaction (RQ5). No significant interactions.
Dimension F p $\omega^{2}$ Sig.
Accuracy 0.66 .876 .000
Lay Relevancy 0.75 .787 .000
Expert Relevancy 0.42 .990 .000
Helpfulness 0.57 .940 .000
Table 43: Strategy $\times$ Model interaction (RQ5). No significant interactions.
Ranking Evaluation.
To address potential positional bias in absolute scoring (gu2024llmasjudge), we conducted a supplementary listwise ranking evaluation. Both judges ranked all 8 strategies for 60 matched sets ($N=120$ rankings total). Table 44 shows average ranks and frequency of best/worst rankings.
Strategy Avg Rank SD #Best #Worst G-Eval
Self-consistency 2.86 1.81 37 2 4.44
Reflexion 3.41 1.74 19 2 4.35
CoT Zero-shot 3.71 1.78 16 1 4.17
Zero-shot 3.92 2.02 16 6 4.35
Role-based 4.14 2.03 14 6 4.32
Few-shot 5.86 1.98 6 22 4.25
CoT Few-shot 6.02 2.01 5 35 4.01
Meta-prompting 6.07 2.05 6 45 4.01
Table 44: Ranking evaluation results (RQ5). Lower rank = better. $N=119$ rankings (combined from 2 judges).
Friedman Test.
Non-parametric repeated-measures test for ranking data: $\chi^{2}(7)=230.3$, $p<.001$. Strategy rankings are significantly different.
G-Eval vs Ranking Correlation.
Spearman correlation between G-Eval average scores (Table 7) and average ranks (Table 44): $\rho=-.857$, $p=.007$. The strong negative correlation (higher G-Eval score $\leftrightarrow$ lower/better rank) validates our pointwise G-Eval evaluation methodology.
Inter-Judge Agreement.
Kendall’s $W=0.72$ for ranking agreement across 60 sets, indicating substantial agreement between judges.